# Graph neural network outputs are almost surely asymptotically constant

### v1

@@ Appendix C Sparse Erdős-Rényi graphs
|$\frac{|\{v\mid\mathrm{IsoTp}_{\ell}(v)=t^{*}\}|}{n}$
is converging to some limit $q_{t^{*}}$ asymptotically almost surely, while by Lemma C.4 we have that:
$\frac{|\operatorname{\mathcal{N}}_{\ell}(\bar{u})|}{n}$
is converging to 00. Hence we can set $q_{t}=q_{t^{*}}$. ∎
We now begin the proof of the sparse case of Theorem 5.1. Recall from the body that we will do this via Lemma 5.8, which requires us to define some Lipschitz controllers on neighbourhood types that approximate a given term $\pi$ relative to a neighbourhood type $t$.
Again we will use $\mathbb{F}$ be the probability distribution of the node features. For each subterm $\pi$ of $\tau$, its reach, denoted $\operatorname{\mathrm{Reach}}(\pi)$ is a natural number defined as follows:
@@
Take a subterm $\pi$ of $\tau$. Let $k$ be the number of free variables in $\pi$. Let $\ell=\operatorname{\mathrm{Reach}}(\pi)$. We now define the controller at $\pi$ for every $t\in\mathsf{IsoTp}_{k,\ell}$. The controller will be of the form:
@@
Note that when $|t|=0$ this is a constant. We can also verify that $\ell$ and $k$ are as in the lemma.
@@
Property C.6.
For every $t\in\mathsf{IsoTp}_{k,\ell}$ and $\epsilon,\theta>0$, there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ when sampling graphs from $\mathrm{ER}(n,p(n))$ we have that for every $k$-tuple of nodes $\bar{u}$ in the graph such that $\mathrm{IsoTp}_{\ell}(\bar{u})=t$, taking the canonical ordering $\operatorname{\mathcal{N}}_{\ell}(\bar{u})=\{s_{1},\ldots,s_{|t|}\}$:
$\left\lVert e_{\pi}^{t}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t|}))-\llbracket\pi(\bar{u})\rrbracket\right\rVert<\epsilon$
@@
    Here $\ell=\operatorname{\mathrm{Reach}}(\pi)=0$. There is only one $(k,0)$ local isomorphism type, namely:
    $t=(\bar{u},\{u_{1},\ldots,u_{k}\})$
    Let $e_{\pi}^{t}(\bar{\bm{a}})\coloneqq{\bm{a}}_{i}$. Note that for any $k$-tuple of nodes $\bar{u}$ in the graph we have:
    $e_{\pi}^{t}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{k}))$ $=\mathrm{H}(s_{i})$
@@
    Then $\ell=\operatorname{\mathrm{Reach}}(\pi)=d$. Note that the random walk embeddings up to length $d$ are entirely determined by the $d$-neighbourhood. Therefore, given local isomorphism type $t$, there is ${\bm{r}}_{t}$ such that:
@@
    for all tuples $\bar{u}$ with local isomorphism type $t$. So set:
@@
    Note that $\ell=\max_{j\leq r}\operatorname{\mathrm{Reach}}(\rho_{i})$. Take $t\in\mathsf{IsoTp}_{k,\ell}$. For each $i\leq r$ let $t_{j}\coloneqq t\restriction\operatorname{\mathrm{Reach}}(\rho_{j})$. Define:
@@
    Now take $\epsilon,\theta>0$. By the induction hypothesis, there is $N$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for every $k$-tuple of nodes $\bar{u}$ in the graph, letting $t=\mathrm{IsoTp}(\bar{u})$, for every $j\leq r$:
    $\left\lVert e_{\rho_{j}}^{t_{j}}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t_{j}|}))-\llbracket\rho_{j}(\bar{u})\rrbracket\right\rVert<\epsilon$
@@
    $\left\lVert e_{\pi}^{t}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t|}))-\llbracket\pi(\bar{u})\rrbracket\right\rVert<L_{F}\epsilon$
@@
    Note that $\ell=\max(\operatorname{\mathrm{Reach}}(\rho),\operatorname{\mathrm{Reach}}(\eta))+1$. Take $t\in\mathsf{IsoTp}_{k,\ell}$.
    When $u_{i}$ has no neighbours according to $t$, define:
@@
    and note that for any $k$-tuple of nodes $\bar{u}$ in the graph such that $\mathrm{IsoTp}_{\ell}(\bar{u})=t$ we have:
    $\llbracket\pi(\bar{u})\rrbracket={\bf 0}=e_{\pi}^{t}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t|}))$
    So suppose that $u_{i}$ has some neighbours according to $t$. For any $\bar{\bm{a}}\in([0,1]^{d})^{|t|}$ and $t^{\prime}\in\operatorname{\mathrm{Par}}(t,u_{i})$, let $\bar{\bm{a}}^{t^{\prime}}$ be the tuple elements of $\bar{\bm{a}}$ corresponding to the elements of $t^{\prime}$.
    Define:
    $e_{\pi}^{t}(\bar{\bm{a}})\coloneqq\frac{f_{\pi}(\bar{\bm{a}},t)}{g_{\pi}(\bar{\bm{a}},t)}$
@@
    $f_{\pi}(\bar{\bm{a}},t)\coloneqq\frac{1}{|\operatorname{\mathrm{Par}}(t,u_{i})|}\sum_{t^{\prime}\in\operatorname{\mathrm{Par}}(t,u_{i})}e_{\rho}^{t^{\prime}}\left(\bar{\bm{a}}^{t^{\prime}}\right)h\left(e_{\eta}^{t^{\prime}}\left(\bar{\bm{a}}^{t^{\prime}}\right)\right)$
@@
    $g_{\pi}(\bar{\bm{a}},t)\coloneqq\frac{1}{|\operatorname{\mathrm{Par}}(t,u_{i})|}\sum_{t^{\prime}\in\operatorname{\mathrm{Par}}(t,u_{i})}h\left(e_{\eta}^{t^{\prime}}\left(\bar{\bm{a}}^{t^{\prime}}\right)\right)$
    Take $\epsilon,\theta>0$. By the induction hypothesis and the fact that $\operatorname{\mathrm{Par}}(t,u_{i})$ is finite, there is $N$ such that for all $n\geq N$ with probability at least $1-\theta$, for every $t^{\prime}\in\operatorname{\mathrm{Par}}(t,u_{i})$ and every $(k+1)$-tuple of nodes $(\bar{u},v)$ such that $\mathrm{IsoTp}_{\ell+1}(\bar{u},v)=t^{\prime}$ we have that both:
    $\left\lVert e_{\rho}^{t^{\prime}}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t^{\prime}|}))-\llbracket\rho(\bar{u},v)\rrbracket\right\rVert<\epsilon$ ($\dagger_{\rho}$)
    and:
    $\left\lVert e_{\eta}^{t^{\prime}}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t^{\prime}|}))-\llbracket\eta(\bar{u},v)\rrbracket\right\rVert<\epsilon$ ($\dagger_{\eta}$)
    Take $n\geq N$. Henceforward we will condition on the event that equation ($\dagger_{\rho}$) and equation ($\dagger_{\eta}$) hold.
    Take any $k$-tuple of nodes $\bar{u}$ in the graph such that $\mathrm{IsoTp}_{\ell}(\bar{u})=t$. It suffices to show, using the definition of the interpretation of the weighted mean operator, that:
    $\left\lVert\frac{\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}\llbracket\rho(\bar{u},v)\rrbracket h(\llbracket\eta(\bar{u},v)\rrbracket)}{\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}h(\llbracket\eta(\bar{u},v)\rrbracket)}-\frac{f_{\pi}^{t}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t|}))}{g_{\pi}^{t}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t|}))}\right\rVert<\iota$
    By Lemma B.6 it suffices to find $\zeta,\xi>0$ which we can make arbitrarily small and constants $\nu,\Omega>0$ such that:
    $\left\lVert\frac{1}{d(u_{i})}\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}\llbracket\rho(\bar{u},v)\rrbracket h(\llbracket\eta(\bar{u},v)\rrbracket)-f_{\pi}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t|}))\right\rVert<\zeta$ (6)
    $\left\lVert\frac{1}{d(u_{i})}\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}h(\llbracket\eta(\bar{u},v)\rrbracket)-g_{\pi}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t|}))\right\rVert<\xi$ (7)
    $\left\lVert\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}\llbracket\rho(\bar{u},v)\rrbracket h(\llbracket\eta(\bar{u},v)\rrbracket)\right\rVert<\Omega d(u_{i})$ (8)
    $\left\lVert\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}h(\llbracket\eta(\bar{u},v)\rrbracket)\right\rVert<\Omega d(u_{i})$ (9)
    $\forall i\leq d\colon\left[\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}h(\llbracket\eta(\bar{u},v)\rrbracket)\right]_{i}>\nu d(u_{i})$ (10)
    Consider equation (7). Since $t$ completely specifies the neighbourhood of $u_{i}$, there is a one-to-one correspondence between the elements of $\operatorname{\mathrm{Par}}(t,u_{i})$ and the elements of $\operatorname{\mathcal{N}}(u_{i})$. Indeed, under the canonical ordering of $\operatorname{\mathcal{N}}_{\ell}(\bar{u})$, every $v\in\operatorname{\mathcal{N}}(u_{i})$ corresponds to the particularisation $t^{\prime}_{v}$ of $t$ at $u_{i}$.
    By equation ($\dagger_{\eta}$) we have that:
    $\left\lVert e_{\eta}^{t^{\prime}_{v}}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t^{\prime}_{v}|}))-\llbracket\eta(\bar{u},v)\rrbracket\right\rVert<\epsilon$
    Hence, by the Lipschitzness of $h$ we have:
    $\left\lVert h\left(e_{\eta}^{t^{\prime}_{v}}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t^{\prime}_{v}|}))\right)-h(\llbracket\eta(\bar{u},v)\rrbracket)\right\rVert<L_{h}\epsilon$
    where $L_{h}$ is the Lipschitz constant of $h$. Therefore:
    $\left\lVert\frac{1}{d(u_{i})}\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}h(\llbracket\eta(\bar{u},v)\rrbracket)-g_{\pi}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t|}))\right\rVert$
    $\leq\frac{1}{d(u_{i})}\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}\left\lVert h\left(e_{\eta}^{t^{\prime}_{v}}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t^{\prime}_{v}|}))\right)-h(\llbracket\eta(\bar{u},v)\rrbracket)\right\rVert$
    $<\frac{1}{d(u_{i})}\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}L_{h}\epsilon$
    $\quad=L_{h}\epsilon$
    The proof of equation (6) is similar.
[…]
    Note that $\ell=0$ and that $\mathsf{IsoTp}_{k,0}$ contains a single element, namely:
    $t=(\bar{u},\{u_{1},\ldots,u_{k}\})$
    Let $\ell^{\prime}=\max(\operatorname{\mathrm{Reach}}(\rho),\operatorname{\mathrm{Reach}}(\eta))$. For any tuple $\bar{u}$ let and $t^{\prime}\in\mathsf{IsoTp}_{k+1,\ell^{\prime}}$:
    $\llbracket t^{\prime}(\bar{u})\rrbracket\coloneq|\{v\mid\mathrm{IsoTp}_{\ell^{\prime}}(\bar{u},v)=t^{\prime}\}|$
    By Corollary C.5, for every $t^{\prime}\in\mathsf{IsoTp}_{k+1,\ell^{\prime}}$ there is $q_{t^{\prime}}\in[0,1]$ such that for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$, for every $k$-tuple $\bar{u}$ we have that:
    Note that we must have:
    $\sum_{t^{\prime}\in\mathsf{IsoTp}_{k+1,\ell^{\prime}}}q_{t^{\prime}}\leq 1$
    Indeed, if this sum were greater than $1$, there would be some finite $S\subseteq\mathsf{IsoTp}_{k+1,\ell^{\prime}}$ such that $\sum_{t^{\prime}\in S}q_{t^{\prime}}>1$, and so by making $\epsilon$ small enough we would have that the proportion of nodes in the graph which satisfy some $t^{\prime}\in S$ is greater than $1$, which is impossible.
    Define the controller:
    $e_{\pi}^{t}({\bm{a}}_{1},\ldots,{\bm{a}}_{k})\coloneqq\frac{f_{\pi}^{t}({\bm{a}}_{1},\ldots,{\bm{a}}_{k})}{g_{\pi}^{t}({\bm{a}}_{1},\ldots,{\bm{a}}_{k})}$
    where:
    $f_{\pi}^{t}({\bm{a}}_{1},\ldots,{\bm{a}}_{k})\coloneqq\sum_{t^{\prime}\in\mathsf{IsoTp}_{k,\ell^{\prime}}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}}\left[\begin{array}[]{c}e_{\rho}^{t^{\prime}}({\bm{a}}_{1},\ldots,{\bm{a}}_{k},{\bm{b}}_{k+1},\ldots,{\bm{b}}_{|t^{\prime}|})\\ h(e_{\eta}^{t^{\prime}}({\bm{a}}_{1},\ldots,{\bm{a}}_{k},{\bm{b}}_{k+1},\ldots,{\bm{b}}_{|t^{\prime}|}))\end{array}\right]q_{t^{\prime}}$
    and:
@@
    $g_{\pi}^{t}({\bm{a}}_{1},\ldots,{\bm{a}}_{k})\coloneqq\sum_{t^{\prime}\in\mathsf{IsoTp}_{k,\ell^{\prime}}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}}\left[h(e_{\eta}^{t^{\prime}}({\bm{a}}_{1},\ldots,{\bm{a}}_{k},{\bm{b}}_{k+1},\ldots,{\bm{b}}_{|t^{\prime}|}))\right]q_{t^{\prime}}$
@@
    Since:
@@
    are bounded, and $\sum_{t^{\prime}\in\mathsf{IsoTp}_{k+1,\ell^{\prime}}}q_{t^{\prime}}$ converges, these sums converges absolutely.
    Now, take $\epsilon,\theta>0$. Take a finite $S\subseteq\mathsf{IsoTp}_{k,\ell^{\prime}}$ such that:
@@
    By Corollary C.5 there is $N_{1}$ such that for all $n\geq N_{1}$ with probability at least $1-\theta$, for all $t^{\prime}\in S$ and every $k$-tuple $\bar{u}$ we have that:

### v2

@@ D.1 Combinatorial tools for the sparse case
|$\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T}_{\ell}(y)\}|}{n}$
converges to some limit $q_{T}$ asymptotically almost surely, while by Lemma D.6 we have that:
$\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})|}{n}$
converges to 00, so the result follows. ∎
D.2 Proving the inductive invariant for the sparse cases, and proving the convergence theorem for these cases
We now begin the proof of Theorem D.1. Recall from the body that we will do this via Lemma 5.7, which requires us to define some Lipschitz controllers on neighbourhood types that approximate a given term $\pi$ relative to a neighbourhood type.
Throughout the following, it will be convenient to assume for every $k$-rooted graph isomorphism type ($T,\bar{u}$) a canonical ordering nodes $T=\{s_{1},\ldots,s_{|T|}\}$ such that:
  • the first $k$ nodes in the ordering are $\bar{u}$ and
  • for any graph $G$, tuple of nodes $\bar{u}$ and $\ell,\ell^{\prime}\in\mathbb{N}$ with $\ell^{\prime}<\ell$ the canonical ordering of $\operatorname{\mathcal{N}}_{\ell^{\prime}}(\bar{u})$ is an initial segment of the canonical ordering of $\operatorname{\mathcal{N}}_{\ell}(\bar{u})$.
Again we will use $\mathbb{F}$ for the probability distribution of the node features. For each subterm $\pi$ of $\tau$, its reach, denoted $\operatorname{\mathrm{Reach}}(\pi)$ is a natural number defined as follows:
@@
Take a subterm $\pi$ of $\tau$. Let $k$ be the number of free variables in $\pi$. We now define the controller at $\pi$ for every $k$-rooted graph $(T,\bar{w})$. The controller will be of the form:
@@
Note that when $|T|=0$ this is a constant.
@@
Property D.7.
Let $\pi$ be any subterm of $\tau$ and let $\ell=\operatorname{\mathrm{Reach}}(\pi)$. Consider sampling a graph $G$ from either the sparse Erdős-Rényi or Barabási-Albert distributions. For every $k$-rooted graph $(T,\bar{w})$ and $\epsilon,\theta>0$, there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for every $k$-tuple of nodes $\bar{u}$ in the graph such that $\operatorname{\mathcal{N}}_{\ell}(\bar{u})\cong(T,\bar{w})$, taking the canonical ordering $\operatorname{\mathcal{N}}_{\ell}(\bar{u})=\{s_{1},\ldots,s_{|T|}\}$:
@@
    Here $\ell=\operatorname{\mathrm{Reach}}(\pi)=0$. Let $e_{\pi}^{T}(\bar{\bm{a}})\coloneqq{\bm{a}}_{i}$, the feature value of the $i$th node in the tuple. Note that for any $k$-tuple of nodes $\bar{u}$ in the graph we have:
    $e_{\pi}^{T}(H_{G}(s_{1}),\ldots,H_{G}(s_{|T|}))$ $=H_{G}(s_{i})$
@@
    Then $\ell=\operatorname{\mathrm{Reach}}(\pi)=d$. Note that the random walk embeddings up to length $d$ are entirely determined by the $d$-neighbourhood. Therefore, given rooted graph $(T,\bar{w})$, there is ${\bm{r}}_{T}$ such that:
@@
    for all tuples $\bar{u}$ such that $\operatorname{\mathcal{N}}_{\ell}(\bar{u})=T$. So set:
@@
    Note that $\ell=\max_{j\leq r}\operatorname{\mathrm{Reach}}(\rho_{i})$. Take a $k$-rooted graph $(T,\bar{w})$. For each $i\leq r$ let $T_{j}\coloneqq\operatorname{\mathcal{N}}^{T}_{\operatorname{\mathrm{Reach}}(\rho_{j})}(\bar{w})$. Define:
@@
    Now take $\epsilon,\theta>0$. By the induction hypothesis, there is $N$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for every $k$-tuple of nodes $\bar{u}$ in the graph, letting $T=\operatorname{\mathcal{N}}_{\ell}(\bar{u})$, for every $j\leq r$:
    $\left\lVert e_{\rho_{j}}^{T_{j}}(H_{G}(s_{1}),\ldots,H_{G}(s_{|T_{j}|}))-\llbracket\rho_{j}(\bar{u})\rrbracket\right\rVert<\epsilon$
@@
    $\left\lVert e_{\pi}^{T}(H_{G}(s_{1}),\ldots,H_{G}(s_{|T|}))-\llbracket\pi(\bar{u})\rrbracket\right\rVert<L_{F}\epsilon$
@@
    In this step we use that the value of a local aggregator is determined by the values of the terms it is aggregating in the local neighbourhood.
    Take a $k$-rooted graph $(T,\bar{w})$, where $k$ is the number of free variables in $\pi$. Note that $\ell=\max(\operatorname{\mathrm{Reach}}(\rho),\operatorname{\mathrm{Reach}}(\eta))+1$.
    When $w_{i}$ has no neighbours in $T$, define:
@@
    and note that for any $k$-tuple of nodes $\bar{u}$ in the graph such that $\operatorname{\mathcal{N}}_{\ell}(\bar{u})=T$ we have:
    $\llbracket\pi(\bar{u})\rrbracket={\bf 0}=e_{\pi}^{T}(H_{G}(s_{1}),\ldots,H_{G}(s_{|T|}))$
    So suppose that $w_{i}$ has some neighbours in $T$. Enumerate the neighbours as:
    $\operatorname{\mathcal{N}}^{T}(w_{i})=\{y_{1},\ldots,y_{r}\}$
    Define the Lipschitz function:
    ${\operatorname*{\mathrm{WMean}}}_{T}({\bm{a}}_{1},\ldots,{\bm{a}}_{r},{\bm{b}}_{1},\ldots,{\bm{b}}_{r})\coloneqq\frac{\sum_{j=0}^{r}{\bm{a}}_{j}h({\bm{b}}_{j})}{\sum_{j=0}^{r}h({\bm{b}}_{j})}$
    Note that whenever $\operatorname{\mathcal{N}}_{\ell}^{G}(\bar{u})\cong(T,\bar{w})$, letting $\operatorname{\mathcal{N}}^{G}(u_{i})=\{v_{1},\ldots,v_{r}\}$ be the enumeration of the neighbourhood of $u_{i}$ given by the isomorphism, we have that:
    $\llbracket\pi(\bar{u})\rrbracket={\operatorname*{\mathrm{WMean}}}_{T}(\llbracket\rho(\bar{u},v_{1})\rrbracket,\ldots,\llbracket\rho(\bar{u},v_{r})\rrbracket,\llbracket\eta(\bar{u},v_{1})\rrbracket,\ldots,\llbracket\eta(\bar{u},v_{r})\rrbracket)$ ($\square$)
    For any $y_{j}\in\operatorname{\mathcal{N}}(w_{i})$, let:
    $T_{j}\coloneqq\operatorname{\mathcal{N}}_{\ell-1}^{T}(\bar{w},y_{j})$
    $T^{\rho}_{j}\coloneqq\operatorname{\mathcal{N}}_{\operatorname{\mathrm{Reach}}(\rho)}^{T}(\bar{w},y_{j})$
    $T^{\eta}_{j}\coloneqq\operatorname{\mathcal{N}}_{\operatorname{\mathrm{Reach}}(\eta)}^{T}(\bar{w},y_{j})$
    Further, for any $\bar{\bm{a}}\in([0,1]^{d})^{|T|}$ let $\bar{\bm{a}}^{\rho}_{j}$ and $\bar{\bm{a}}^{\eta}_{j}$ be the tuple of elements of $\bar{\bm{a}}$ corresponding to the nodes of $T^{\rho}_{j}$ and $T^{\eta}_{j}$ respectively.
    Let:
    $e_{\pi}^{T}(\bar{\bm{a}})\coloneqq{\operatorname*{\mathrm{WMean}}}_{T}\left(e_{\rho}^{T^{\rho}_{1}}(\bar{\bm{a}}^{\rho}_{1}),\ldots,e_{\rho}^{T^{\rho}_{r}}(\bar{\bm{a}}^{\rho}_{r}),e_{\eta}^{T^{\eta}_{1}}(\bar{\bm{a}}^{\eta}_{1}),\ldots,e_{\eta}^{T^{\eta}_{r}}(\bar{\bm{a}}^{\eta}_{r}),\right)$
    Take $\epsilon,\theta>0$. Applying the induction hypothesis to the term $\rho(\bar{u},v)$ and to $\eta(\bar{u},v)$, which have $\operatorname{\mathrm{Reach}}$ at most $\ell-1$, and using the fact that $\operatorname{\mathcal{N}}^{T}(w_{i})$ is finite, there is $N$ such that for all $n\geq N$ with probability at least $1-\theta$, for every $y_{j}\in\operatorname{\mathcal{N}}^{T}(w_{i})$ and every $(k+1)$-tuple of nodes $(\bar{u},v)$ such that $\operatorname{\mathcal{N}}^{G}_{\ell-1}(\bar{u},v)\cong(T_{j},\bar{w},y_{j})$ we have that both:
    $\left\lVert e_{\rho}^{T^{\rho}_{j}}(H_{G}(s_{1}),\ldots,H_{G}(s_{|T^{\rho}_{j}|}))-\llbracket\rho(\bar{u},v)\rrbracket\right\rVert<\epsilon$
    and:
    $\left\lVert e_{\eta}^{T^{\eta}_{j}}(H_{G}(s_{1}),\ldots,H_{G}(s_{|T^{\eta}_{j}|}))-\llbracket\eta(\bar{u},v)\rrbracket\right\rVert<\epsilon$
[…]
    $\left\lVert e_{\pi}^{T}(H_{G}(s_{1}),\ldots,H_{G}(s_{|T|}))-\llbracket\pi(\bar{u})\rrbracket\right\rVert<L_{T}\epsilon$
    where $L_{T}$ is the Lipschitz constant of ${\operatorname*{\mathrm{WMean}}}_{T}$.
  • Inductive step for the GCN aggregator $\pi={\textsc{GCN}}_{v\in\operatorname{\mathcal{N}}(u_{i})}\rho$.
    This step is very similar to the previous, where we again use the fact that the value of a local aggregator is determined by the values of the terms it is aggregating in the local neighbourhood. The only difference is that we now have a different Lipschitz function:
    ${{\textsc{GCN}}}_{T}({\bm{a}}_{1},\ldots,{\bm{a}}_{r})\coloneqq\sum_{j=1}^{r}\frac{1}{\left|\operatorname{\mathcal{N}}^{T}(u_{i})\right|\left|\operatorname{\mathcal{N}}^{T}(y_{j})\right|}{\bm{a}}_{j}$
    The rest of the proof is as before.
  • Inductive step for global weighted mean $\pi=\sum_{v}\rho\star h(\eta)$.
    Note that $\ell=0$. Let $\ell^{\prime}=\max(\operatorname{\mathrm{Reach}}(\rho),\operatorname{\mathrm{Reach}}(\eta))$. For any tuple $\bar{u}$ and $(k+1)$-rooted graph $(T^{\prime},\bar{u},v)$ let:
    $\llbracket T^{\prime}(\bar{u})\rrbracket\coloneq|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{w},y)\cong T^{\prime}\}|$
    By the parameterized Weak Local Convergence result, Theorem D.3, for every such $T^{\prime}$ there is $q_{T^{\prime}}\in[0,1]$ such that for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$, for every $k$-tuple $\bar{u}$ we have that:
    Let $\mathcal{T}_{k+1}$ be the set of isomorphism types of $(k+1)$-rooted graphs. Note that:
    $\sum_{T^{\prime}\in\mathcal{T}_{k+1}}q_{T^{\prime}}\leq 1$
    Indeed, if this sum were greater than $1$, there would be some finite set $S\subseteq\mathcal{T}_{k+1}$ such that $\sum_{T^{\prime}\in S}q_{T^{\prime}}>1$, and so by making $\epsilon$ small enough we would have that the proportion of nodes in the graph which satisfy some $T^{\prime}\in S$ is greater than $1$, which is impossible.
    Consider any $(k+1)$-rooted graph $(T^{\prime},\bar{w},y)$. In order to define the controller, we need to identify the nodes in the canonical enumeration corresponding to $\operatorname{\mathcal{N}}_{\ell^{\prime}-1}^{T^{\prime}}(\bar{u})$. Let:
    $r_{T^{\prime}}\coloneqq\left|\operatorname{\mathcal{N}}_{\ell^{\prime}-1}^{T^{\prime}}(\bar{u})\right|$
    Given tuples of nodes $\bar{\bm{a}}=({\bm{a}}_{1},\ldots,{\bm{a}}_{r_{T^{\prime}}})$ and $\bar{\bm{b}}=({\bm{b}}_{r_{T^{\prime}}+1},\ldots,{\bm{b}}_{|T^{\prime}|})$, let $\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}})$ be the $|T^{\prime}|$-tuple of vectors obtained by assigning the ${\bm{a}}_{i}$’s to the nodes in $\operatorname{\mathcal{N}}_{\ell^{\prime}-1}^{T^{\prime}}(\bar{u})$ and the ${\bm{b}}_{i}$’s to the remaining nodes in $T^{\prime}$, according to the canonical enumeration of $T^{\prime}$.
    Define the controller:
    $e_{\pi}^{T}(\bar{\bm{a}})\coloneqq\frac{f_{\pi}^{T}(\bar{\bm{a}})}{g_{\pi}^{T}(\bar{\bm{a}})}$
@@
    $f_{\pi}^{T}(\bar{\bm{a}})\coloneqq\sum_{T^{\prime}\in\mathcal{T}_{k+1}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}}))\cdot h(e_{\eta}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}})))\right]\cdot q_{T^{\prime}}$
@@
    $g_{\pi}^{T}({\bm{a}}_{1},\ldots,{\bm{a}}_{k})\coloneqq\sum_{T^{\prime}\in\mathcal{T}_{k+1}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[h(e_{\eta}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}})))\right]\cdot q_{T^{\prime}}$
@@
    This completes the definition of the controller. We now present the proof that it is correct.
    We first argue that the sum in the definition is well-defined. Since:
@@
    are bounded, and $\sum_{T^{\prime}\in\mathcal{T}_{k+1}}q_{T^{\prime}}$ converges, these sums converges absolutely.
    Now, take $\epsilon,\theta>0$. Take a finite $S\subseteq\mathcal{T}_{k+1}$ such that:
@@
    The guarantee given when we applied the parameterized version of Weak Local Convergence, Theorem D.3, is that there is $N_{1}$ such that for all $n\geq N_{1}$ with probability at least $1-\theta$, for all $T^{\prime}\in S$ and every $k$-tuple $\bar{u}$ we have that: